Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below

Sales (12,600 units × $30 per unit) $ 378,000
Variable expenses 189,000
Contribution margin 189,000
Fixed expenses 211,500
Net operating loss $ (22,500 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,800 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $86,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $38,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 0.60 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,900?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $52,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,900 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,900)?
Answer #1

PEM Inc

  1. Computation of CM ratio and break-even point in unit sales and dollar sales:

CM ratio = (contribution margin/sales) x100

Contribution margin = $189,000

Sales = $378,000

CM Ratio = (189,000/378,000) x 100 = 50%

Break-even point in unit sales –

Break-even point in unit sales = fixed cost/contribution margin per unit

Contribution margin per unit = $189,000/12,600 units = $15

Fixed cost = $211,500

Break-even point in units sales = 211,500/15 = 14,100 units

Break-even point in dollar sales –

Break-even point in dollar sales = fixed cost/CM ratio

= $211,500/50% = $423,000

  1. Determination of the increase or decrease in the company’s monthly net operating income –

increase in fixed cost (advertising budget) = $6,800,

revised fixed cost = 211,500 + 6,800 = $218,300

Increase in monthly sales = $86,000

Revised sales = $378,000 + $86,000 = $464,000

Contribution margin = 50% of sales = $464,000 x 50% = $232,000

Net operating income = contribution margin – fixed cost

= $232,000 - $218,300 = $13,700

Original net operating income = ($22,500)

Revised net operating income = $13,700

Change = 13,700 – (22,500) = 13,700 + 22,500 = $36,200

Hence, net operating income increased by $36,200.

  1. Determination of revised net operating income:

Sales price reduced by 10%, = $30 -10% $30 = $27

Revised fixed cost (with increased advertising expenditure) = $211,500 + $38,000 = $249,500

Revised unit sales = 2 x 12,600 = 25,200 units

Revised contribution margin = revised sales – revisedvariable cost

Revised sales = $27 x 25,200 units = $680,400

Revised variable cost = $15 x 25,200 units = $378,000

Revised CM= 680,400 – 378,000 = $302,400

Revised CM per unit = $302,400/25,200 units= $12 per unit

Revised net income = Revised CM – Revised fixed expenses

= $302,400 - $249,500 = $52,900

Hence, revised net income = $52,900

  1. Determination of number of units to be sold to reach monthly target of $4,900:

Revised variable cost = $15 + $0.60 = $15.60 per unit

Revised CM per unit = $30 - $15.60 = $14.40

Fixed expenses = $211,500

Target income = $4,900

Desired sales units = (fixed expenses + target profit)/CM per unit

= (211,500 + 4,900)/14.40 = 15,028 units

Hence, the company needs to produce and sell 15,028 units to reach the target profit of $4,900.

  1. A. New CM ratio and break-even point in unit sales and dollars:

Break-even point in unit sales -

Revised variable cost = $15 - $3 = $12 per unit

New CM per unit = $30 - $12 = $18

New CM ratio = new CM per unit/sales price per unit

= (18/30) x 100 = 60%

Revised Fixed expenses = 211,500 + 52,000 = $263,500

Break-even point in unit sales = fixed expenses/CM per unit

= $263,500/$18 = 14,639 units

Break-even point in dollar sales = fixed expenses/CM ratio

= 263,500/60% = $439,167

5B      

Contribution margin income statements

Automation

No Automation

Per Unit

Total

Percentage

Per Unit

Total

Percentage

Sales units

20,900

20,900

Sales

$30

$627,000

100%

$30

$627,000

100%

Variable cost

$12

$250,800

40%

$15

$313,500

50%

Contribution margin

$18

$376,200

60%

$15

$313,500

50%

Fixed expenses

$263,500

42.00%

$211,500

33.70%

Net Income

$112,700

18.00%

$102,000

16.30%

5c.

Yes, we recommend the company adopt automation.

Automation resulted in increase in net income by $10,700 (112,700 – 102,000). Though automation increases fixed expenses by $52,000 it causes a reduction of $3 per unit in variable costs and increase in net income by $10,700. Hence, automation is recommended.